OpenStudy (anonymous):

2/4+1/5=something over 20 ?

OpenStudy (anonymous):

2/4 + 1/5 You must get a common denominator in order to add. The lowest common denominator is 20. Therefore, 10/20 + 4/20 = 14/20. However, this is reducible to 7/10.

OpenStudy (anonymous):

If you don't understand the common denominator part, think of it this way. You can only add fractions if the have the same bottom (denominator). So, multiply the fraction by whatever it takes the denominators to be the same. For 2/4, you would need to multiply the whole thing by 5/5, which gets you 10/20. Now, if you multiply it by 5/5, it's really the same as multiplying it by 1. Make sense? 5/5 = 1. Also, you end up with the same overall "Value" of the number. 2/4 = 10/20. So really, all your doing it changing the look of the fractions so you can add them. You're not really changing their overall values. Hopefully this clarifies it up a bit.